### v1

# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

Abstract
Humans can continuously acquire new skills and knowledge by exploiting existing ones for improved learning, without forgetting them. Similarly, ‘continual learning’ in machine learning aims to learn new information while preserving the previously acquired knowledge. Existing research often overlooks the nature of human learning, where tasks are interleaved due to human choice or environmental constraints. So, almost never do humans master one task before switching to the next. To investigate to what extent human-like learning can benefit the learner, we propose a method that interleaves tasks based on their ‘learning progress’ and energy consumption. From a machine learning perspective, our approach can be seen as a multi-task learning system that balances learning performance with energy constraints while mimicking ecologically realistic human task learning. To assess the validity of our approach, we consider a robot learning setting in simulation, where the robot learns the effect of its actions in different contexts. The conducted experiments show that our proposed method achieves better performance than sequential task learning and reduces energy consumption for learning the tasks.
Index Terms:
interleaved learning, multi-task learning, human-like learning, energy budget, learning progress, intrinsic motivation.
I Introduction
Humans possess a remarkable ability to learn continuously, seamlessly integrating new information while retaining previously acquired knowledge. This dynamic learning process is characterized by the constant interleaving of tasks and subjects; we often switch tasks in response to a dynamic balance of environmental and cognitive constraints, adjusting behavior in real-time when continuing a task becomes less efficient or more mentally costly [1]. This fluid interplay of switching between different activities not only enhances memory consolidation but also adapts us to dynamic environments by shifting attention toward differences between activities and forcing us to be more mentally engaged [2]. In the realm of artificial intelligence, mimicking this continuous and interleaved learning paradigm remains a significant challenge.
Traditional machine learning approaches often rely on isolated task learning, which optimizes performance for a specific task but neglects knowledge transfer from related tasks, thereby limiting generalization and adaptability across domains [3]. While multi-task learning attempts to improve generalization by training models on multiple tasks simultaneously [4], continual learning tackles sequential task acquisition by learning tasks one after another [5, 6]. Yet, both approaches still fall short of replicating the flexible, interleaved learning strategies observed in humans.
The human brain has evolved to optimize energy efficiency, favoring strategies that maximize information transmission per unit energy [7]. Through mechanisms such as synaptic pruning and neural plasticity [8], it optimizes its neural pathways, ensuring efficient cognitive function while reducing its metabolic costs [9, 10]. Along with the brain’s drive toward efficient information processing, we also get inspiration from learning progress (LP), which has been proposed as an intrinsic motivation mechanism that guides exploration based on the rate of improvement in performance. Schmidhuber [11, 12] introduced LP as a measure of how quickly an agent reduces its prediction error, encouraging the agent to focus on experiences that yield the highest learning gains. This aligns well with biological principles, as the brain appears to allocate information encoding resources preferentially to stimuli that are neither too simple nor too complex, maximizing informational value relative to effort [13, 14].
In this paper, we propose a novel approach that integrates interleaved multi-task learning with energy-modulated learning progress. By designing a system that not only switches between tasks in an interleaved fashion but also adjusts its learning dynamics based on energy considerations, we aim to create a more human-like learning model. Our method is evaluated in simulated robotic environments focused on effect prediction tasks—specifically, having the robot predict the outcome of the interacted objects. The experimental results demonstrate that suitable task switching leads to an interleaved learning regime, which improves learning performance across tasks. Additionally, introduction of energy-modulation to the learning progress-based task selection results in reduced energy consumption without a significant drop in learning efficiency. This indicates that our approach not only aligns more closely with human learning patterns but also offers practical benefits in terms of efficiency and resource management. By bridging the gap between human continuous learning and machine learning methodologies, this work contributes to the development of more adaptable and efficient artificial intelligence systems. Our findings suggest that embracing interleaved learning and energy modulation can significantly enhance the capability of machines to learn in a manner that mirrors human cognition, opening avenues for future research in continuous and sustainable learning models.
II Related Work
II-A Multi-task Learning
Multi-task learning (MTL) [15] is a machine learning paradigm in which several tasks are learned jointly so that knowledge contained in one task can help improve the performance of the others. Unlike transfer learning—where the emphasis is on improving one target task, MTL treats all tasks as equally important and seeks to boost them collectively. MTL is also distinct from multi-label learning and multi-output regression because in MTL, the tasks typically do not share the same data records (i.e., each task has its own dataset). The key questions pertaining to MTL are (1) when to use MTL to share knowledge, (2) what to share (features, instances, or parameters, and (3) how, e.g., by feature learning, clustering, or parameter sharing, where ’how’ is the most commonly studied in the literature [4].
In feature learning, it is common to map the original features [16, 17, 18, 19, 20, 21] to common space or choosing a subset of the original task features [22, 23]. One of the notable architectures in the feature learning category is the Cross-Stitch Networks[24]. It dynamically combines feature maps from parallel task-specific networks, by a cross-stitch unit that learns a linear combination of the activation maps from the previous layer, allowing the model to adaptively determine which features to share based on the input data. This work has the closest architecture to our previous work [25], which laid the groundwork for this study. However, Cross-Stitch Networks differ in key aspects from our work They are typically initialized with task-specific single-task networks, finetuned on the respective task, and then combined using cross-stitch units to perform joint training. In contrast, in our model, all tasks are trained from scratch in a unified architecture without any pretraining or fine-tuning. In addition, after introducing the cross-stitch units, in their approach, all the tasks are trained jointly, using the same input data, where we only train one task per training iteration, which is chosen by our task arbitration mechanism. Other than the feature-based models, parameter-based models are also developed for MTL problems such as low-rank methods [26, 27] suggesting that if tasks are related, the matrix of model parameters often has low rank, task clustering methods [28, 29, 30] that hypothesize tasks form separate clusters of similarity, with each group sharing parameters. Other parameter-based approaches such as task relation [31, 32] and parameter matrix decomposition [33, 34] are likewise covered in the literature.
An open question in MTL, other than learning and skill sharing, is how to balance the contributions of individual tasks that may vary widely in complexity and data distribution. Recent research has addressed this through adaptive loss weighting and gradient balancing. Kendall et al. [35] introduced an uncertainty-based framework that scales each task’s loss according to its inherent noise, thereby directing the learning process toward more reliable signals. Complementing this, Chen et al. [36] proposed GradNorm, a method that normalizes the gradients of different tasks to ensure that no single task dominates the optimization process.
In addition, Sener and Koltun [37] reframed MTL as a multi-objective optimization problem, seeking Pareto-optimal solutions that effectively navigate trade-offs between conflicting task objectives. MTL approaches adopting mixture-of-expert architectures are also proposed, such as in [38], where task contributions are dynamically weighted and shared representations are exploited. However, despite the effectiveness of these approaches in balancing learning objectives and modeling task relations, they typically assume simultaneous or sequential task training schedules. In contrast, our work departs from this by introducing a dynamic task arbitration mechanism that selects a single task at each training step based on learning progress and energy consumption. This enables an interleaved training regime that more closely resembles human learning patterns, which we describe in the following section.
II-B Interleaved Learning in Humans
Interleaved learning is a cognitive strategy in which learners alternate among diverse topics or problem types within a single study session, rather than concentrating on one subject exclusively, thereby fostering flexible thinking and long-term retention. Empirical research in cognitive psychology provides compelling evidence for the benefits of interleaved practice. For example, studies have demonstrated that interleaving mathematics problems enhance learners’ ability to distinguish between problem types, leading to improved problem-solving skills [39]. Similarly, research by Kornell and Bjork [40] indicates that interleaved learning facilitates the formation of more flexible and integrated representations of concepts, enabling learners to better apply acquired knowledge in new contexts.
Neuroimaging studies further show that interleaving practice drives increased frontal–parietal activity and heightened motor cortex excitability along with the reduced retrieval time of information, thus leading to better long-term retention and efficient retrieval compared to the blocked practice [41]. These interleaved practice effects result from a phenomenon called contextual interference (CI) [42], which is one of the “desirable difficulties” [43] that suggest introducing a challenge during the learning can lead to an improvement in long term retention [44]. In a more recent study by Rohrer et al.[45], math problems are shuffled so that the problems belonging to the same kind are not consecutively solved by the students, requiring students to come up with a proper strategy while solving the problem, similar to the real world situations. The experiments showed that interleaved practice produced higher scores compared to blocked practice, in a final test given on both day 1 and day 30 (delayed test), showcasing protection against forgetting. In addition, the study suggests that interleaved learning benefits do not lessen over time; quite the opposite, they may increase over time. Similarly, in [46] showed that undergraduate physics students who practiced interleaved problem sets demonstrated improved memory and problem-solving skills compared to those who used blocked practice. These findings highlight the potential of interleaved learning as a powerful strategy for optimizing human learning in both educational and real-world settings.
In machine learning, however, interleaved learning received very little attention. Recently, Mayo et al. [47] investigated the interleaving on a multi-task learning problem and discussed that rather than designing mechanisms to prevent forgetting, such as external memory [48] or regularization of weights [49, 50], we should focus on designing learning systems and schedules that embrace the natural and resilient mechanisms of human learning, where forgetting is not a failure but a feature that coexists with the capacity for efficient recovery i.e., relearning savings [51] and long-term knowledge retention. Even without mechanisms to prevent forgetting, standard neural networks show memory retention effects [47], similar to humans, when tasks are interleaved.
II-C Learning Progress: An Intrinsic Motivation
Intrinsic motivation (IM) refers to the drive to engage in activities for their inherent satisfaction, rather than for external rewards [52]. In the context of AI and robotics, intrinsic motivation enables agents to exhibit behaviors such as curiosity [53, 54], novelty [55, 56], and surprise [57], similar to the motivations observed in humans [58, 59, 60]. Another IM that is often used in robotic and computational approaches is the prediction progress, also known as learning progress (LP) [61], such that improvement of predictions over time leads to the generation of rewards. It is usually calculated by comparing the predictor’s error before and after it is updated, using the same sensorimotor context [11]. There are numerous applications of LP in machine learning problems such as exploration guiding in reinforcement learning [62], region selection [63, 64]. Colas et. al used LP in their multi-goal RL model (CURIOUS) as an IM signal to select which goal module to practice and replay, prioritizing those with the highest absolute LP to drive efficient and adaptive curriculum learning. In the same vein, in our previous work [25] that forms the conceptual basis of the current study, we used LP as a signal to autonomously arbitrate task selection. Unlike prior works that use LP solely for guiding exploration, goal or region selection, this paper applies LP at the task level and integrates it with energy consumption to enable efficient, human-like interleaved multi-task learning.
II-D Energy Conservation of the Brain
Energy management in the human brain is not only about powering neural activity—it also reflects a sophisticated system for conserving energy that influences our behavior [65]. Despite the brain’s relatively small size, it consumes a disproportionate share of the body’s energy, mainly to support neural signaling through the maintenance of ion gradients and synaptic transmission [66]. Over time, the brain has evolved mechanisms such as synaptic pruning and circuit rewiring, where frequently used pathways are strengthened, and rarely used connections are eliminated. This dynamic reorganization minimizes unnecessary activity, leading to what is often described as neural efficiency [67, 68]. In addition to the brain’s physiological mechanisms for preserving energy, research in cognitive neuroscience and psychology has also shown that people systematically avoid tasks perceived as highly demanding, requiring more effort or energy, making the associated rewards seem less valuable,e which is often explained by the term effort discounting [69]. For example, Kool et al. [70] demonstrated that individuals tend to choose less cognitively demanding tasks when given the option, reflecting an inherent preference for minimizing mental effort. Similarly, Westbrook and Braver [71] provided evidence that cognitive effort carries a subjective cost, influencing decision-making processes. This concept is further elaborated in the opportunity cost model proposed by Kurzban et al. [72], which suggests that the perceived cost of expending cognitive energy plays a central role in our choices, leading us to opt for tasks that require minimal resource expenditure. In this study, we are inspired by the findings from neural and cognitive science and propose an energy-efficient task selection mechanism in addition to learning-based task arbitration.
II-E Effect Prediction and its Applications
Effect prediction tasks in robotics involve endowing agents with the capability to anticipate the outcomes of their actions, similar to the humans [73], forming an internal model of the environment that can be used for planning and control [74]. This predictive capacity—often referred to as a forward model [75] is fundamental to intelligent behavior. By simulating the effects of different actions, a robot can evaluate potential strategies [76] before committing to a course of action, thereby enhancing safety, efficiency, and adaptability.
At the core of effect prediction is the idea that the robot learns to map its actions to subsequent states. Early approaches focused on building explicit physical models [77, 78]; however, recent advances have shifted towards learning these models directly from high-dimensional sensory data using deep learning techniques. For example, Watter et al. [79] introduced a latent dynamics model that transforms raw sensory inputs, such as images, into a compact, low-dimensional latent space where the dynamics of the environment are more predictable. In this latent space, the model approximates the effect of actions as locally linear transformations, which can be leveraged for short horizon planning. Another work by Agrawal et al. [80] focuses on self-supervised learning of intuitive physics. Their system trains a robot to predict the outcome of simple interactions—like poking objects—without explicit supervision. Over time, the model learns to infer the underlying physical properties of objects, which is crucial for manipulating unfamiliar items. Furthermore, effect prediction is central to model-based reinforcement learning (RL). In model-based RL, the agent uses its learned internal model to simulate future trajectories, allowing it to plan over longer time horizons while reducing the need for extensive trial-and-error interactions with the real environment. The “World Models” approach by Ha and Schmidhuber [81] exemplifies this trend by learning a compact representation of the world that supports internal simulation and planning. This internal simulation capability enables robots to achieve higher sample efficiency and improved performance, particularly in complex or resource-constrained settings.
In robotic applications, effect prediction tasks play a pivotal role in different problems, such as navigation, manipulation, and autonomous planning. In manipulation, [82] emphasizes the role of object- and relation-centric representations in improving push effect prediction, showing how modeling inter-object dynamics allows robots to infer and control the consequences of their actions in cluttered scenes. This low-level understanding of physical interactions feeds directly into symbolic and goal-directed planning, as seen in [83], who propose the Multi-Object Graph Affordance Network (MOGAN) to model compound object affordances and enables planning with multi-object interactions. Building on this, Ahmetoglu et al. [84] demonstrate how effect prediction supports symbolic manipulation planning by learning object and relational predicates, providing a bridge between physical action and abstract reasoning. Similarly, in another work [85], they introduce symbolic attentive layers, which allow the robot to ground predictive object symbols in sensory data, enabling generalizable planning across diverse tasks. While these works primarily focus on manipulation and planning, the principles extend to navigation as well; Aktas et al. [86] show that predicting the effects of partial action executions enables more flexible, multi-step planning, which is essential for navigating dynamic or partially observable environments. However, these studies did not use intrinsic motivation signals during exploration and relied on the observations from random interactions with the environment.
By incorporating effect prediction into learning frameworks, as demonstrated in the works mentioned above, robotics can achieve a level of adaptability and efficiency akin to human behavior. Just as humans use their internal models to simulate potential outcomes and opt for actions that minimize risk and energy expenditure, robots equipped with predictive models can choose actions that optimize performance while conserving resources. This alignment with biological principles not only advances the field of robotics but also inspires the development of more sustainable, energy-aware artificial intelligence systems.
III Problem Statement
The problem of interest is to answer how a dynamical task arbitration mechanism for data sampling and learning can lead to better overall learning and skill transfer in multitask learning. The multitask learning framework we consider follows human learning, where the learner has to decide which task to engage in and when to disengage, effectively allowing the emergence of arbitrary interleaved learning regimes.
Specifically, in the context of supervised regression problems, we design an online multi-task learning model equipped with a task arbitration mechanism that enables the agent to decide, at each training step, which task to focus on for data sampling and learning. Formally, given $T$ tasks where task $t\in\{1,2,\dots,T\}$, in each training step, the task arbitration mechanism chooses a task $t^{*}\in T$ to be learned by the multi-task learning agent. Next, the agent interacts with the corresponding task environment $\mathcal{E}_{t^{*}}$ and collects new data
$\mathcal{D}_{t^{*}}=\{(x_{i}^{(t^{*})},y_{i}^{(t^{*})})\}_{i=1}^{N_{t^{*}}},$
where $x_{i}^{(t^{*})}$ represents the $i$-th input for task $t^{*}$, $y_{i}^{(t^{*})}$ is the corresponding target and $N_{t^{*}}$ is the number of samples of task $t^{*}$. This approach contrasts with the standard offline setup, where a complete dataset is available prior to the beginning of training. Once the agent gathers the $\mathcal{D}_{t^{*}}$, it then updates its specific parameter set $\theta^{(t^{*})}$ along with the $\theta^{s}$ which is the parameter set shared among the tasks. This dynamic process continues until the end of the training procedure. The predictive function for each task $t\in T$ is given by
$f_{t}\bigl{(}x^{(t)};\theta^{s},\theta^{(t)}\bigr{)}.$
Each task is associated with a loss function $\mathcal{L}_{t}$, which quantifies the error between the predicted outputs and the true labels. The objective is to find parameters $\theta^{s},\{\theta^{(t)}\}_{t=1}^{T}$ that minimize the total loss across all tasks:
$\min_{\theta^{s},\{\theta^{(t)}\}}\sum_{t=1}^{T}\mathcal{L}_{t}\left(\left\{\left(f_{t}(x_{i}^{(t)};\theta^{s},\theta^{(t)}),\,y_{i}^{(t)}\right)\right\}_{i=1}^{N_{t}}\right).$
Coming to the effect prediction tasks in this work, an agent is required to learn a total of $T$ distinct effect prediction tasks by establishing a mapping from the state-action space to the effect space. For each task $t\in\{1,2,\ldots,T\}$, the state is represented as $x^{(t)}\in\mathbb{R}^{n_{t}}$ and the action as $a^{(t)}\in\mathbb{R}^{m_{t}}$, where $n_{t}$ and $m_{t}$ denote the dimensions of the state and action spaces, respectively. The corresponding effect, which is the outcome of executing action $a^{(t)}$ in state $x^{(t)}$, is denoted by $e^{(t)}\in\mathbb{R}^{k_{t}}$. The goal is to dynamically arbitrate data sampling and learning among the tasks so that for each task $t$ action-effect prediction functions, $f_{t}:\mathbb{R}^{n_{t}}\times\mathbb{R}^{m_{t}}\rightarrow\mathbb{R}^{k_{t}}$, are learned. After learning, it is expected that for any state-action pair $(x^{(t)},a^{(t)})$ the predicted effect $\tilde{e}^{(t)}=f_{t}(x^{(t)},a^{(t)})$ closely approximates the actual effect $e^{(t)}$. To formalize this further, we can define the data set an agent is able to collect with
$\mathcal{D}_{t}=\{(x_{i}^{(t)},a_{i}^{(t)},e_{i}^{(t)})\}_{i=1}^{N_{t}}$
where each tuple provides a sample of the state, action, and resulting effect.
IV Methodology
Figure 1: Overview of the proposed multi-task effect prediction architecture and an example training iteration for $t^{*}=2$ where total number of tasks $T=3$. The model features shared and task-specific components, with only the selected task being trained at each step based on the arbitration mechanism.
In this section, we detail the architecture and methodologies employed in our proposed interleaved multi-task learning framework with energy-modulated learning progress. Our goal is to realize human-like learning by interleaving tasks and optimizing energy consumption through a specialized neural network architecture and task selection mechanism.
IV-A Neural Network Architecture
Our model is based on an encoder-decoder network architecture designed to handle multiple-task learning while allowing for inter-task information sharing. The key components of the architecture are task-specific input (state) and action projection layers, task-specific encoders as well as a shared encoder and a shared attention module, and finally task specific output (effect) decoders. We detail these components next.
  1. 1.
    State Projection Layer. For each task $t$ in the set of tasks $T$, the state input $x^{(t)}\in\mathbb{R}^{n_{t}}$ is first passed through a task-specific projection function $P_{state}^{(t)}\colon\mathbb{R}^{n_{t}}\to\mathbb{R}^{d_{s}}$ where $n_{t}$ is the input dimensionality of task $t$ and $d_{s}$ is the dimensionality of the projected state input. This layer maps the input into a fixed-dimensional space to ensure consistency across different tasks. Mathematically, the projected input is given by:
    $x^{\prime(t)}=P_{state}^{(t)}(x^{(t)}),$ (1)
    where $x^{\prime(t)}\in\mathbb{R}^{d_{s}}$. This step ensures that inputs from different tasks are compatible for subsequent shared processing. For the implementation of this layer, either a linear or nonlinear function, depending on the specific problem, can be chosen. However, since the primary objective of these layers is to map inputs of varying dimensions onto a uniform dimensional space, a linear projection function is implemented in our work, as often done in the literature [87, 88].
  2. 2.
    Action Projection Layer. Since the action vectors $a^{(t)}\in\mathbb{R}^{m_{t}}$ for each task $t$ can vary in dimensionality, similar to the state projection layers, we project them to a fixed-dimensional space to maintain consistency across tasks. Each action vector is passed through a task-specific projection function $P_{action}^{(t)}\colon\mathbb{R}^{m_{t}}\to\mathbb{R}^{d_{a}}$:
    $a^{\prime(t)}=P_{action}^{(t)}(a^{(t)}),$ (2)
    where $a^{\prime(t)}\in\mathbb{R}^{d_{a}}$ is the projected action vector for task $t$ and $d_{a}$ is the fixed dimensionality of the projected action space. This projection ensures that actions from different tasks are compatible for further processing. Similar to the (1), we choose to have a linear projection layer for the action vectors.
  3. 3.
    Shared State Encoder. We incorporate a shared encoder in our neural network with the motivation that it will extract general features that are beneficial across multiple tasks, thus promoting knowledge sharing while reducing redundancy, leading to more efficient learning. The projected input $x^{\prime(t)}$ is then fed into the a shared encoder $F\colon\mathbb{R}^{d_{s}}\to\mathbb{R}^{d_{h}}$, to extract low-level features common across different tasks:
    $h=F(x^{\prime(t)}),$ (3)
    where $h\in\mathbb{R}^{d_{h}}$ is the encoded representation, and $d_{h}$ is the dimensionality of the shared encoder output.
  4. 4.
    Task State Encoders. Following the shared encoder, further, we integrate task-specific encoders aiming to capture state knowledge unique to each task that the shared encoder cannot extract. We hypothesize that each task benefits from specialized features, enhancing its performance during training. Hence, the $h$ is forwarded through all task-specific encoders in order to generate task-specific latent representations:
    $r^{(t)}=f^{(t)}(h),\quad\forall t\in T$ (4)
    where $r^{(t)}\in\mathbb{R}^{d_{r}}$ is the task-specific latent representation, $f^{(t)}\colon\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{r}}$ is the task-specific encoder function of $t$ and $d_{r}$ is the dimensionality of the task specific latent space. Note that during the training of task $t^{*}$, for all $f^{(t)}$ where $t\neq t^{*}$ are frozen, which means gradient propagation is not carried out for them.
  5. 5.
    Shared Attention Module. To facilitate inter-task communication and let the network focus on the current training task, we utilize a multi-head attention (MHA) mechanism [89]. Instead of having one attention function, where $Q,K,V\in\mathbb{R}^{d_{model}}$, MHA layer computes attention across $H$ parallel heads that each head projects $Q,K,V$ into a lower-dimensional space and performs scaled dot-product attention. Let $d_{k}$ be the dimensionality per head, and let $W_{i}^{Q},W_{i}^{K},W_{i}^{V}\in\mathbb{R}^{d_{model}\times d_{k}}$ be the projection matrices for head $i\in\{1,\dots,H\}$. Then, following [89] we have:
    $Q_{i}=Q\,W_{i}^{Q},\quad K_{i}=K\,W_{i}^{K},\quad V_{i}=V\,W_{i}^{V},$
    $\mathrm{head}_{i}=\mathrm{softmax}\!\Bigl{(}\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}\Bigr{)}\,V_{i}\in\mathbb{R}^{d_{k}}.$
    The outputs from all heads are concatenated and passed through an output projection matrix to form the final attention output, denoted here as $A$, as follows:
    $A=\left[\mathrm{head}_{1}\;:\;\mathrm{head}_{2}\;:\;\dots\;:\;\mathrm{head}_{H}\right]W^{O}\in d_{model},$
[…]
    • Beforehand the attention function, we concatenate an additional flag bit to each task-specific latent representation, $r^{(t)}$. Specifically, for each task $t$, we concatenate a training flag $\delta^{(t)}$ indicating whether task $t$ is currently being trained $(\delta^{(t)}=1)$ or not $(\delta^{(t)}=0)$. The concatenated vector for each task is:
      $z^{(t)}=[r^{(t)}\;:\;\delta^{(t)}]\in\mathbb{R}^{d_{r}+1}$ (5)
      From now on, the $z^{(t)}$ here is the final representation of task $t$.
    • After gathering the $z^{(t)}$ for each task, they are stacked to form a matrix $Z$:
      $Z=\begin{bmatrix}z^{(1)}\\ z^{(2)}\\ \vdots\\ z^{(T)}\end{bmatrix}\in\mathbb{R}^{T\times(d_{r}+1)}$ (6)
      where $T$ is the number of tasks, $d_{r}$ is the fixed dimensionality of task-specific latent representations. Extra dimension accounts for the training flag.
    • Once we obtain the matrix $Z$, the $Q,K$ and $V$ values are gathered as below:
      Query ($Q$)
      We choose the final representation of the current training task as the query since we want the attention mechanism to be driven by the perspective of the currently active task:
      $Q=z^{(t^{*})}\in\mathbb{R}^{(d_{r}+1)},$
      where $d_{r}+1$ is the embedding dimension, or $d_{model}$.
      Key ($K$) and Value ($V$)
      We use $Z$ so that the active training task can attend to representations from all tasks (including itself):
      $K=V=Z\in\mathbb{R}^{T\times(d_{r}+1)},$
    Using these $Q,K,V$ values, we feed them into our shared MHA module and get the attention output $A$ for the task $t=t^{*}$:
    $A^{(t)}=SharedMHA(Q,K,V),$ (7)
    where $A^{(t)}\in\mathbb{R}^{(d_{r}+1)}$. Because the query is restricted to the representation of the active task, $z^{(t^{*})}$, the network focuses on how the other tasks’ representations (contained in $Z$) can inform task $t^{*}$. Moreover, the flags within each row of $Z$ indicate whether a given input originates from the actively trained task ($\delta^{t=t^{*}}=1$) or from another task ($\delta^{t\neq t^{*}}=0$), further guiding the attention toward balancing shared and task-specific information.
  6. 6.
    Task Effect Decoders. Same as the task-specific sub-encoder $f^{(t)}$, for each task $t$; there is a decoder module $g^{(t)}$ designed for predicting the resulted output (effect) $e^{(t)}$, after the executed action $a^{(t)}$ while the environment is in the state $x^{(t)}$. For task $t=t^{*}$, the output from the shared MHA mechanism $A^{(t)}$ along with the projected action vector $a^{\prime(t)}$ from (2) is given to the task-specific decoder of the training task only, contrary to the (4), as follows:
    $\tilde{e}^{(t)}=g^{(t)}(A^{(t)},a^{\prime(t)}).$ (8)
    Here $\tilde{e}^{(t)}\in\mathbb{R}^{k_{t}}$ and $k_{t}$ is the output dimension of task $t$. By integrating contextual information from the shared attention mechanism with the projected action vector, the decoder maps the state information to the corresponding effect.
IV-B Task Arbitration
To investigate the effects of interleaving on multi-task learning, we propose a task selection strategy based on the dynamic assessment of each task’s learning progress and, additionally, its associated energy consumption. Unlike traditional multi-task learning approaches that train tasks in a fixed schedule or isolate them entirely, our method continuously evaluates the tasks’ performance trends and selects which task to train next based on these evaluations. By doing so, we simulate the human tendency to switch between tasks in response to learning progress and energetic constraints. Below, the interleaved task selection algorithm is explained in detail:
Algorithm 1 Interleaved Multi-Task Learning
0:  Number of tasks $T$, number of epochs $K$, exploration rate $\epsilon=0.1$
  for epoch $=1$ to $K$ do
     Compute score $s^{(t)}$ for each task $t\in\{1,2,\ldots,T\}$
[…]
     Generate a random number $r$ uniformly from $[0,1]$
     if $r<\epsilon$ then
        $t^{*}\leftarrow$ Random task from $\{1,2,\ldots,T\}\setminus\{t_{\text{winner}}\}$
     else
        $t^{*}\leftarrow t_{\text{winner}}$
     end if
     Perform training on the selected task $t^{*}$
  end for
  1. 1.
    Learning Progress (LP) Based Task Selection Similar to LP definition in[11], in our work LP is assessed by examining the recent evolution of a task’s error signal. For a given task $t$, let $E^{(t)}_{i}$ be the error at time step $i$. We track the error over the last $L$ training steps (e.g., $L=5$) to compute the slope of the error curve. Consider the set of errors $\{E^{(t)}_{i-L+1},E^{(t)}_{i-L+2},\ldots,E^{(t)}_{i}\}$. We fit a linear model to these points to estimate the slope $\beta^{(t)}$. If $\beta^{(t)}<0$, the error is decreasing, and the task is making positive learning progress. If $\beta^{(t)}\geq 0$, it suggests that the error has plateaued or is increasing, implying no recent improvement. Under the LP-based interleaving scheme, we compute the learning progress $LP^{(t)}$ for each task $t$. We then select the task with the highest LP (most improving) to be trained next. In other words, tasks that are currently showing rapid improvement receive more training time, while tasks that have stalled or regressed receive less immediate attention. Formally, if $\beta^{(t)}$ is the slope for task $t$, then we choose $t^{*}$ in two steps:
    1. (a)
      ${LP^{(t)}}=\begin{cases}\lvert\beta^{(t)}\rvert,&{\text{if }}\ \beta^{(t)}<0\\ {0,}&{\text{otherwise.}}\end{cases}$
    2. (b)
[…]
    This ensures that at any given moment, the model focuses on tasks where additional practice can yield meaningful gains, rather than spending time on tasks that have stagnated.
  2. 2.
    Energy Consumption (EC) Modulated Learning Progress (LP) Based Task Selection. While learning progress alone can guide the model toward tasks showing improvement, it does not consider the computational or energetic cost associated with training each task. In scenarios where energy efficiency is a concern, we incorporate energy consumption into the interleaving strategy. We define the neurons’ activation “energy” consumption $EC^{(t)}$ for task $t$ as the cumulative energy used over the last $L$ training steps. This could be measured in terms of computational operations, memory usage, or any proxy for energy expenditure. Lower $EC^{(t)}$ values mean the task has recently been trained efficiently, while higher values indicate that the task has been relatively costly in terms of energy. However, using EC only as the interleaving guidance, the network will only choose the tasks that produce low activation energy without taking the learning performance of the tasks into consideration, which can lead to worsened overall multi-task learning performance (see Figure 7a). To blend both learning performance and energy considerations, we employ the following combined score for each task $t$:
    $s^{(t)}=\exp{(k\cdot LP^{(t)})}/EC^{(t)},$ (9)
    where:
    • $LP^{(t)}$ is the learning progress defined above.
    • $EC^{(t)}$ is recent energy consumption for the task.
    • $k$ is a positive constant that controls the sensitivity of the combined score to energy consumption ($EC^{(t)}$).
    Note that, to combine LP and EC in a balanced manner, we first normalize both values between $[0,1]$ before computing the combined score. The $s^{(t)}$ can be interpreted as follows:
    • When $k$ approaches to zero, $\exp{(k\cdot LP^{(t)})}\approx 1$, thus the score is primarily influenced by $EC^{(t)}$, even if the task has strong LP. This discourages spending too many resources on a single, costly task.
    • When $k$ is high, $\exp{(k\cdot LP^{(t)})}$ becomes dominant, and thus EC becomes negligible.
    Under the EC-modulated LP-based interleaving strategy, we compute combined scores for each task $t$ as given above. The next task to train is the one with the highest combined score:
[…]
    This ensures a balance between quick gains in task performance and maintaining overall energy efficiency. Tasks with high learning progress but also low recent energy costs are favored, while tasks that have become too energy-intensive or are not showing improvement receive less immediate focus. By integrating these interleaving strategies, we encourage a more dynamic, human-like learning schedule. The LP-based method shifts focus toward tasks that are currently “improvable,” while the EC-modulated LP method adds a layer of energy awareness, guiding the interleaving process toward both efficiency and sustained improvement. An overview of the task arbitration mechanism used in this work is presented in Algorithm 1, where a non-greedy selection strategy is adopted by giving other tasks a chance to be selected, with an exploration rate $\epsilon$, thereby reducing over-reliance on the highest-scoring task.
IV-C Simulation Environment
In our study, we use a simulated tabletop environment to evaluate multi-task learning through effect prediction tasks. The simulation is implemented using PyBullet engine [90] and utilizes a UR10 robotic arm equipped with a Franka Panda end effector. The tabletop is a constrained environment designed to replicate real-world conditions for object interaction tasks. A variety of objects with differing shapes, sizes, and physical properties are placed on the table for the robot to interact with during the tasks. The objects used in this study include spheres, cubes, cylinders, and square prisms. Cylinder and square prism can be placed either horizontally or vertically, resulting in a total of six distinct object types.
IV-D Effect Prediction Tasks
In this work, we define three action-effect prediction tasks performed in the simulated environment explained above. Each task requires the model to predict the resulting state of one or more objects after a specific action is executed. The effect prediction tasks are designed as follows:
Push task. A single object is randomly placed on the table. The robot applies a “pushing” action with its end effector to the object’s center of mass (CoM), with an angle chosen between $[0,180]$ degrees. The state of the object includes its Cartesian position (x, y, z) and orientation, represented using the sine and cosine of each Euler angle (x, y, z axes), resulting in a 9-dimensional state representation:
$\mathbf{s}=[x,y,z,\sin(\phi_{x}),\cos(\phi_{x}),$
$\sin(\phi_{y}),\cos(\phi_{y}),\sin(\phi_{z}),\cos(\phi_{z})].$
The action is encoded using the sine and cosine of the push angle, combined with a one-hot vector indicating the object’s id:
$\mathbf{a}=[\sin(\theta),\cos(\theta),\text{onehot}(o)].$
Hit Task. The second task is a dynamic variation of the Push task. The setup is the same, but the pushing action is applied with twice the velocity, resulting in more chaotic and less predictable object behavior. The state representation and action encoding remain identical to the Push task.
Stack Task. In the final task, the goal is to place one object on top of another, forming a possible stable stack. This task involves two objects: a moving object and a target object. As a result, the state includes the positional and orientation information for both objects, doubling the state information compared to the Push and Hit tasks.
$\mathbf{s}=[\mathbf{s}_{m},\mathbf{s}_{t}],$
where $\mathbf{s}_{m}$ and $\mathbf{s}_{t}$ are the 9-dimensional states of the moving and target objects, respectively. The action is represented by concatenating the one-hot encodings of both the moving object and the target object:
$\mathbf{a}=[\text{onehot}(o_{m}),\text{onehot}(o_{t})].$
V Experiments and Results
To evaluate the effectiveness of our proposed method, which utilizes EWLP-based task selection within an interleaved multi-task learning framework, we perform a series of experiments across different aspects. The central aim of these experiments is to investigate how intelligent task interleaving influences learning efficiency, knowledge transfer, and final task performance in comparison to other approaches. We begin by comparing INTER-LP to several baselines designed to isolate the contributions of individual components. These include a single-task learning setup (SINGLE), a multi-task model with random task selection (INTER-RAND), and a blocked learning setup where tasks are trained sequentially in fixed permutations (BLOCK). In addition to these, we conduct a set of experiments on our energy-aware variant, INTER-LPE, which extends INTER-LP by incorporating energy consumption into the task arbitration mechanism. These experiments are designed to explore how varying the energy sensitivity coefficient $k$ affects the trade-off between prediction performance and computational cost. By analyzing different values of $k$, we demonstrate the flexibility of INTER-LPE in adapting to resource constraints while maintaining competitive learning outcomes. Since we maintain a consistent model architecture across all settings, we ensure that any observed performance differences are attributable to task arbitration strategies rather than architectural complexity or capacity. The experiments are structured to evaluate overall learning performance across training regimes, investigating task-wise learning dynamics to examine how individual tasks benefit from shared training and interleaving, assessing robustness to network complexity by comparing models across low, medium, and high parameter counts, demonstrating the advantages of interleaved learning over blocked training in mitigating catastrophic forgetting, and conducting ablation studies to reveal the individual and joint contributions of architectural components such as the shared attention mechanism and task-specific flag bits. Together, these evaluations provide a comprehensive picture of how and why EWLP-guided interleaved learning can outperform traditional scheduling strategies in multi-task environments.
V-A Baselines
To demonstrate the advantages of our interleaved multitask learning method with LP-based task selection (INTER-LP), we compare it against the following baselines:
  • SINGLE: To isolate the impact of shared representation (multi-task learning), our first baseline is a single-task learning model where each task has its own dedicated network for training. The proposed network in Figure 1 is adapted for single-task learning where each task has its own state projection, encoder, action projection, attention, and decoder modules without any shared parameters between tasks. While it is possible to implement single-task learning using a simpler vanilla MLP for each task, we intentionally adopt the same multi-task learning architecture used for our proposed method. This ensures that the comparison is fair, with both multitask and single-task setups having identical model structures, allowing the differences in performance to be attributed solely to the learning strategy rather than architectural discrepancies.
  • INTER-RAND: The second baseline uses the shared architecture used in our proposed method, without any modification but removes the LP-based task selection mechanism. Instead, tasks are chosen randomly at each training iteration, with each task having an equal chance of being selected. This baseline serves to highlight the impact of our task selection (interleaving) strategy by comparing it against a scenario where the interleaving is uniformly random. This baseline helps isolate the effectiveness of the scheduling algorithm from the benefits of shared parametrization.
  • BLOCK: The third and final baseline is implemented using the same shared network architecture but trains tasks in dedicated, uninterrupted blocks rather than interleaving them. To account for any potential effects caused by the order in which tasks are learned, we evaluate all six possible permutations of the three tasks (denoted as BLOCK-xyz, where the xyz represents the order of task training with their corresponding ids). In our case, we have three tasks, namely Push (0), Hit (1), and Stack (2), and train for a total of 3000 epochs. Thus, for example, “BLOCK-012” means that the Push task is trained alone for the first 1000 epochs, followed by a Hit task for the next 1000, and a Stack task for the remaining 1000. This baseline enables us to assess whether the interleaved approach offers improved learning over all possible blocked learning schedules.
V-B Training Settings
All models are implemented using the PyTorch [91] framework, using the same set of hyper-parameters across all baseline methods and our proposed model: a learning rate of $0.001$, a batch size of $100$, a hidden dimension of $4$, and training for $3000$ epochs. Each encoder and decoder module, whether shared or task-specific, consists of two fully connected layers, with ReLU used as the activation function. Optimization is performed using the AdamW optimizer [92], with the AMSGrad variant enabled and default weight decay settings.
Although the SINGLE baseline ends up with a higher number of parameters per task compared to all other models, including ours, we maintain a consistent network architecture across all settings. This design choice ensures that performance differences arise primarily from differences in task interleaving strategies, rather than architectural complexity.
V-C Overall Learning Performance
The initial evaluation of the proposed model compares it to the SINGLE and INTER-RAND baselines. Experiments were conducted using 10 different random seeds, and the average task performance results are presented here. As illustrated in Figure 2, the proposed model INTER-LP reduces prediction error faster than the other two baselines. This indicates that the LP-based task selection with inter-task skill transfer produces an interleaved task schedule that leads to superior overall performance compared random selection and independent learning with no skill transfer. Furthermore, it is important to note that not only does interleaving enhance performance, but simultaneously learning multiple tasks is also beneficial, as tasks can mutually support each other and facilitate positive information transfer. This is evidenced by the INTER-RAND baseline showing improvement compared to the single-task baseline.
Figure 2: Overall performance of baselines is shown for isolated task learning (SINGLE), proposed architecture but random task selection (INTER-RAND) and proposed architecture with LP-based task selection model (INTER-LP. As can be seen proposed model (INTER-LP) converges faster than the baselines. Even though the random task selection is slightly better than the single task learning, it could not surpass learning with LP-based task selection
Task-wise Analysis. To better understand the learning dynamics, we examine the performance of individual tasks throughout the learning process. As illustrated in Figure 3, the performance on both the Push and Stack tasks improves when all tasks are learned simultaneously, with a marked enhancement observed when using the proposed model. Notably, the Stack task exhibits a significant performance gap between our proposed model, INTER-LP and the other two baselines. This discrepancy likely arises from the Stack task’s greater complexity compared to the other tasks, demonstrating that single-task learning is insufficient for handling such complexity. Hit task, on the other hand, receives a brief early learning improvement during iterations of $300-1000$ and does not show as large gains as the other two tasks. Although the INTER-RAND baseline utilizes a multi-task learning approach, it results in little to no improvement when interleaving is not performed suitably.
Figure 3: Task-specific model performance is shown. It can be seen that LP-based method surpasses the other baselines on both push and stack tasks, with a higher significance in the latter. However, for the hit task, the performance is not improved considering the other two baselines.
V-D Network Complexity
Figure 4: Model performances with different neural network complexity levels. Even though at high parameter regimes, all models perform similarly, the LP-based learning method (INTER-LP) becomes superior at low or mid-parameter regimes, where suitable task switching for appropriate skill transfer among tasks becomes crucial.
To analyze the impact of network complexity on performance, we evaluate baselines (SINGLE and INTER-RAND) and our proposed model (INTER-LP) as follows: Each model is trained with three levels of network complexity: Low ($\sim 800$ parameters), Medium ($\sim 2000$ parameters), and High ($\sim 5200$ parameters). The results, as shown in Figure 4, illustrate how performance evolves across different resource levels.
Low Complexity. At the lowest network complexity, INTER-LP achieves a significantly lower MAE compared to the two baselines, indicating its ability to efficiently utilize limited resources. This result highlights the effectiveness of our learning progress-based interleaving strategy in scenarios where computational resources are constrained. In contrast, SINGLE and INTER-RAND exhibit higher prediction errors, suggesting that they are less capable of optimizing task performance when resources are limited.
Medium Complexity. As the network size increases, the performance gap between INTER-LP and the two baselines begins to narrow. While INTER-LP still outperforms both SINGLE and INTER-RAND, the relative improvements are less pronounced compared to the low-complexity scenario. This indicates that the baselines start benefiting from the additional parameters, but INTER-LP continues to leverage its task-interleaving strategy to maintain superior performance.
High Complexity. At the highest level of network complexity, the difference between INTER-LP and the baselines is further reduced, with all three models achieving relatively similar levels of performance. This result suggests that as the network’s capacity grows, both SINGLE and INTER-RAND can partially compensate for their limitations in interleaving strategy by relying on the increased capacity for task representation. However, INTER-LP remains competitive, showcasing better performance across all network complexities.
V-E Interleaved vs. Blocked Learning
Figure 5: Interleaved learning versus blocked learning in all possible orders (titles showing the training order of the tasks). No matter the order in which the tasks are trained, the interleaved learning method always surpasses the blocked learning method, which shows that interleaved learning prevents catastrophic forgetting of learned information by switching the context in periods, thus being able to recall faster.
We conducted a detailed comparison of interleaved learning and blocked learning in a multi-task learning network across the three tasks (Push, Hit, and Stack). The interleaved learning mode, implemented in our model INTER-LP, alternates between tasks during training, while the blocked learning mode, implemented in the BLOCKED model, learns each task sequentially in its entirety before moving to the next.
Given the three tasks, there are six possible training orders (e.g., Push $\rightarrow$ Hit $\rightarrow$ Stack or Push $\rightarrow$ Stack $\rightarrow$ Hit, etc.). To see whether any order can perform similar or bettered that interleaved learning, we trained the BLOCKED model across all six task orders, while the INTER-LP model remained unchanged across these configurations, as it does not depend on the task order. Figure 5 shows the performance of the INTER-LP model and the BLOCKED model for each of the six task orders. It can be seen that the proposed model, INTER-LP consistently outperforms the BLOCKED model across all task orders. This consistent performance demonstrates the robustness of interleaved learning, as it mitigates the catastrophic forgetting observed in the Blocked model. Specifically, in blocked learning, the model exhibits a marked decrease in performance on earlier tasks as it progresses through subsequent tasks. For instance, when training starts with a Push task in the BLOCKED model, its performance on a Push task significantly degrades by the time Hit and Stack are learned. Additionally, the variability in performance across different task orders for the BLOCKED model highlights its sensitivity to the order of training. This variability further emphasizes the limitations of blocked learning in scenarios where task interdependence or order effects play a critical role.
V-F EWLP Based Task Selection Performance
Since we hold that the human brain also considers computational cost, in this section, we explore discounting the learning progress of the tasks by the neural cost they incur for task switching. We look at different discount levels and compare the learning performances. Figure 6 presents the overall learning performance of our proposed method (INTER-LP), its energy-weighted variants (INTER-LPE-K1 and INTER-LPE-K1.2), and the baseline models (SINGLE and INTER-RAND), in terms of prediction performance. As expected INTER-LP shows the lowest mean absolute error, with a marked performance margin during the early and mid stages. The energy-weighted variants, INTER-LPE-K1 and INTER-LPE-K1.2, also outperform both SINGLE and INTER-RAND baselines, indicating that incorporating energy awareness does not nullify the gains of our architecture. In fact, INTER-LPE-K1.2 shows performance nearly on par with INTER-LP, suggesting that slightly increasing the weight of learning progress relative to energy cost (via the sensitivity constant $k$) can further enhance training efficiency. The SINGLE model exhibits the weakest performance overall, reaffirming that task isolation limits skill learning efficiency. Meanwhile, INTER-RAND shows modest improvements over SINGLE, highlighting that even random task interleaving can offer some benefit, yet remains clearly inferior to structured LP-based scheduling. Taken together, these results confirm that both learning progress and energy-aware arbitration strategies provide meaningful gains in training efficiency and final task performance, with INTER-LP and INTER-LPE variants converging faster and reaching lower prediction errors than all baselines.
Figure 6: Comparison of baselines and LP-based model to the EWLP-based models where the sensitivity constant is set to $k=1$ and $k=1.2$, respectively.
To further evaluate the effect of EWLP-based task scheduling, we investigate the performance of the INTER-LPE model under varying values of the energy sensitivity coefficient $k$, and compare it to INTER-LP and SINGLE baselines. As shown in Figure 7a, which reports the MAE exactly at epoch 1000, INTER-LP achieves the lowest prediction error overall, confirming the strong effectiveness of LP-based task arbitration. Then, INTER-LPE-K1.2 closely follows, demonstrating that modest energy modulation can preserve much of the learning performance. As $k$ decreases, the prediction performance gradually degrades, indicating that overly prioritizing energy conservation (i.e., smaller $k$) can hinder learning effectiveness. On the other hand, Figure 7b shows total energy consumption measured in terms of cumulative neuron activations. Here, we observe the opposite trend: energy usage decreases steadily with smaller $k$, with INTER-LPE-K0.4 consuming the least energy overall, almost the same with the SINGLE baseline. While INTER-LP achieves the best predictive performance, it is also the most energy-intensive. Notably, INTER-LPE-K1.2 and K1 provide a favorable trade-off—offering significant reductions in energy consumption with only a small sacrifice in prediction accuracy. This tunable behavior highlights the flexibility of the INTER-LPE framework, enabling agents to balance performance and energy efficiency according to task demands and environmental constraints.
(a) Training performances of the models.
(b) Energy consumption of models.
Figure 7: Effect of varying the energy sensitivity coefficient $k$ in the INTER-LPE model on (a) prediction performance and (b) total energy consumption. Higher values of $k$ prioritize learning progress over energy efficiency, leading to lower prediction error but increased energy usage, while lower $k$ values reduce energy consumption at the cost of predictive accuracy.
V-G Ablation Experiments
V-G1 Necessity of Attention and the Flag Bit
Figure 8: Contributions of the shared attention and flag bit from the proposed multi-task learning architecture to the model performances. Neither attention nor flag alone can surpass the convergence speed of our proposed model, which combines them together.
To evaluate the contributions of the attention layer and the flag bit in our proposed model INTER-LP, we conducted an ablation study by training four variations of the model under different configurations. Each model was trained using 10 different random seeds to ensure the robustness of the results. Below, we describe the four settings and provide a detailed analysis of their performance:
Ablation: none. This is our proposed model, which includes a shared attention layer taking the matrix $Z$ from (6), which is formed using the task representations along with their corresponding flag bits. The flag bits serve to differentiate tasks, while the attention layer dynamically learns task interactions.
Ablation: flag. This variation retains the shared attention layer but removes the flag bits. Task representations are combined and passed to the shared attention layer without explicit differentiation using flags.
Ablation: attention. In this model, the flag bits are concatenated to the task representations, the same as the original matrix $Z$ (6), but the shared attention layer is removed from the network.
Ablation: attention + flag. This configuration removes both the shared attention layer and the flag bits. Only the row-wise concatenated task representations are passed to the task decoder of the current training task.
The results of the experiments that are shown in Figure 8 demonstrate the following:
Ablation: attention + flag. This model exhibited the worst performance, consistently underperforming throughout training compared to the other configurations. The absence of both the attention layer and flag bit limits the ability of the model to utilize shared information across tasks or explicitly distinguish between tasks. This indicates that both mechanisms are crucial for enabling effective multi-task learning.
Ablation: attention. Incorporating only the flag bit improved performance compared to the previously ablated model. The flag bit enables task differentiation by augmenting task representations with explicit task-specific information. However, the lack of a shared attention mechanism reduces the model’s ability to capture dynamic task relationships, resulting in suboptimal performance compared to models with attention.
Ablation: flag. This configuration, which retains the shared attention layer but removes the flag bit, outperformed the previous two models. The attention mechanism facilitates dynamic interaction across tasks, allowing the model to better leverage shared information, even in the absence of explicit task differentiation through flag bits.
Ablation: none (ours). The original configuration, which combines both the shared attention layer and the flag bit, achieved the best performance across all training epochs. This result highlights the complementary nature of the two components: the attention layer dynamically learns shared task interactions, while the flag bit provides explicit task differentiation. Together, these mechanisms enable the model to outperform all other configurations consistently, starting from the early stages of training.
VI Conclusion
This paper introduced a biologically inspired interleaved multi-task learning framework that selects tasks dynamically based on learning progress and energy consumption. Motivated by how humans interleave tasks and regulate cognitive effort, our method, INTER-LP, prioritizes tasks that exhibit positive learning progress, while the energy-aware extension, INTER-LPE, further modulates selection by discounting the LP values, leading to a preference for the task with lower energy consumption. The model follows a shared encoder-decoder structure equipped with an attention mechanism that can learn multiple tasks with dynamic task switching. Our approach is evaluated on a set of action-effect prediction tasks in a simulated robotic environment.
The experimental results demonstrate that LP-based task arbitration not only improves overall learning performance and convergence speed compared to baseline strategies like random interleaving, single-task learning, and blocked training, but also facilitates beneficial knowledge transfer across tasks. Moreover, the inclusion of energy-awareness in the task selection process enables a tunable trade-off between learning accuracy and resource efficiency, which is especially valuable in energy-constrained settings. Our ablation studies further validate the complementary roles of the attention mechanism and task flag bits in enhancing multi-task learning dynamics.
While the model is tested in a controlled simulation environment with a limited number of tasks, the results point to broader implications for continuous and sustainable learning systems. By drawing from principles of intrinsic motivation and neural efficiency, the proposed framework contributes a novel perspective on how task interleaving and energy-awareness can be adopted in artificial agents. Future work may explore extensions to real-world robotic platforms, integration with reinforcement learning settings, or scaling to larger task sets with more diverse dynamics.